OpenStudy (cheyennecoffelt):

100K=.85m3 150K=1.27m3 200K=1.7m3 250K=2.12m3 300K=2.54m3 350K=2.97m3 400K=3.39m3 What is the relationship between the temperature and the volume? what will the approximate volume be at 1273.15K?

OpenStudy (anonymous):

Do you see a pattern between temperature and volume?

OpenStudy (cheyennecoffelt):

no, i tried adding, subtracting, all that, my parents even tried to help but we don't know.

OpenStudy (anonymous):

The temperature from 100K to 150K to 200K to 250K. Is it increasing or decreasing? K is Kelvins.

OpenStudy (cheyennecoffelt):

it's increasing

OpenStudy (anonymous):

Correct! Now for lets do volume. The units for volume is \[m^3\] it can be any (length) to the power of 3 Volume goes from 0.85 m^3 to 1.27 m^3 to 1.7m^3 Is it increasing or decreasing?

OpenStudy (cheyennecoffelt):

The volume is also increasing

OpenStudy (anonymous):

Yes, so if temperatures increases does volume increases or decreases?

OpenStudy (cheyennecoffelt):

As the temperature increases, the volume also increases.

OpenStudy (anonymous):

Yes that is the relationship between volume and temperature.

OpenStudy (anonymous):

For the second one, You see that temperature increases by 50K each time right? For volume 0.85m^3 to 1.27 m^3 to 1.7m^3 to 2.12m^3 to 2.54m^3 I would subtract it to see if there is a pattern. 1.27-0.85=0.42 1.7-1.27=0.43 2.12-1.7=0.42 2.54-2.12=0.42 So lets say the pattern is that every time temperature increases by 50K, the volume increases by 0.42m^3 as well.

OpenStudy (anonymous):

1273.15K-400K=873.15K But for every 50K, volume increases by 0.42m^3 So you divide 873.15K/50=? The number you get from dividing; you would then multiply it by 0.42m^3 to get your answer.

OpenStudy (cheyennecoffelt):

okay thank you sooo much, now i can move on! :D

OpenStudy (anonymous):

No problem! Glad you understood Thanks for being very responsive and cooperative.